a car is travelling at the rate of 120 mph while plane is flying at the rate of 600 mph how much longer will it take the car to travel the distance of 400 miles than the plane takes to travel a distance of 1800 miles
Answer by Alan3354(69443)

a car is travelling at the rate of 120 mph while plane is flying at the rate of 600 mph how much longer will it take the car to travel the distance of 400 miles than the plane takes to travel a distance of 1800 miles
===============
d = r*t
t = d/r
----
difference = (400/120) - (1800/600) hours